OpenStudy (anonymous):

Which of the following are measurements of the sides of a right triangle? 101, 99, 20 28, 26, 12 17, 14, 6 none of the above

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

your job is to check whether \[a^2+b^2=c^2\] so for example, is \[20^2+99^2=101^2\]?

OpenStudy (anonymous):

Do you know what a right triangle means?

OpenStudy (misty1212):

if it is, say "yes" if not, say "no"

OpenStudy (anonymous):

Okay give me a moment :)

OpenStudy (anonymous):

no

OpenStudy (misty1212):

last one is not either http://www.wolframalpha.com/input/?i=6%5E2%2B14%5E2%3D17%5E2

OpenStudy (anonymous):

so, it's the second one?

OpenStudy (misty1212):

no i don't think so http://www.wolframalpha.com/input/?i=12%5E2%2B26%5E2%3D28%5E2

OpenStudy (misty1212):

i would go with "none of the above"

OpenStudy (anonymous):

Okay, thank you. :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]